OpenStudy (anonymous):

Please help me. The price of a video game is $32.99 before tax. Andre bought the video game for 20% off. He then paid 8% sales tax on the discounted price. Part A: How much did Andre pay in sales tax? Show all work and steps in your solution. Part B: What is the total amount that Andre paid for the video game? Show all work and steps in your solution.

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@zepdrix can you please help me?

OpenStudy (anonymous):

Do you understand what 20 percent off means in terms of mathematical operations?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

okay so then take 8 percent of that and that is the tax

OpenStudy (anonymous):

to find part b add the tax to the sale price and that will be the total amount he paid

OpenStudy (anonymous):

Wait can you start over and explain it on A

OpenStudy (anonymous):

so 20 percent of 32.99 means \[.2\times32.99\] then subtract that value from 32.99 \[32.99-(.2\times32.99)\] this is the sales price.

OpenStudy (anonymous):

26.392

OpenStudy (anonymous):

@563blackghost can you help me finish this?

563blackghost (563blackghost):

one sec let me finish helping someone^^

OpenStudy (anonymous):

K Thanks (:

563blackghost (563blackghost):

So you would first multiply 32.99 by 0.2 since 0.2 is the decimal form of the percentage of 20%.... \[\Large32.99 \times 0.2 = 6.59\] Now you would get 6.59 which is the what would be taken off the original price... \[\Large32.99 - 6.59 = 26.40\] 26.40 is the price of the video game with the discount now we need to add in the sales tax...So we would multiply 26.40 with 0.08 since 0.08 is decimal form of the percent of 8%... \[\Large26.40 \times 0.08 = 2.11\] 2.11 is the sales tax....Now we would add the discount total price with the sales tax.... \[\Large26.40 + 2.11 = ?\] This will give you the total amount... Sorry I rushed a lil its cause I gtg....

563blackghost (563blackghost):

Correct ^^

563blackghost (563blackghost):

So A would be 2.11 since it is the sales tax....And B would be the total $28.51

OpenStudy (anonymous):

oh... I see. Thank you so much @563blackghost (:

563blackghost (563blackghost):

\(\Large{No~Problem!}\)

OpenStudy (anonymous):

i have two more questions. can i ask you them?

563blackghost (563blackghost):

Im sorry but I must go... :(

OpenStudy (anonymous):

oh its ok. Thanks anyway. Have a good day (:

563blackghost (563blackghost):

Happy I was of help ^^ byeeeee...you have a good day too xD Good Luck on your work